OpenStudy (anonymous):

Translate the following situation into an equation. Do not solve. “Kenny spent one hour more than Lucas studying for their Chemistry test. Andrea spent one hour less than twice the hours Lucas spent studying for the test. If together they spent fourteen hours studying, how many hours did Lucas study?” please show work!

OpenStudy (anonymous):

Hours for Lucas =x Hours for Kenny=x+1 hours for Andrea=2x-1 Total =14 hours So, x+x+1+2x+1=14 then, 4x+2=14

OpenStudy (anonymous):

I dont think is 3

OpenStudy (anonymous):

if u check:: 3+ (3+1) + (2*3)-1)= 7 + (6-1) = 7 + 5 = 12 12 not equqal to 14

OpenStudy (anonymous):

keebler the question u wrote seem not right either ,can u check I solve it not right

OpenStudy (anonymous):

made sure u write correct?

OpenStudy (anonymous):

my answer it decimal ,not even

OpenStudy (anonymous):

i figured out the answer its (x)+(x+1)+(2x-1)=14

OpenStudy (anonymous):

it says do not slove!

OpenStudy (anonymous):

but Nancy u are correct!

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

the frist Iwrote te eqution the same with u, but I solve ,I not sure it correct

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

YeH, that's just the unsimplified answer I wrote.